# The Commensal Real-time ASKAP Fast Transient incoherent-sum survey

### v1

@@ 7.1 FRB host galaxies
Over the course of the survey, we undertook a series of studies investigating the properties of FRB hosts. Bhandari et al. (2020b) and Heintz et al. (2020) conducted the first demographic studies of FRB host galaxies, finding them to be largely late-type star forming galaxies. Bhandari et al. (2022) extended the analysis to include a larger sample of FRBs and compared the properties of the hosts of apparently one-off FRBs with those of repeating sources. The sample size is insufficient to distinguish any difference between the repeating and non-repeating sources. Gordon et al. (2023) undertook detailed modelling of host galaxy spectral energy distributions to assess the star formation histories of nearly all our FRB hosts identified prior to the beginning of 2022. From these studies it is now clear that the host galaxies of one-off bursts are markedly different from that of the first repeating FRB 20121102A which was found to originate in a metal-poor dwarf galaxy Tendulkar et al. (2017).
@@ 7.2 Burst-emission physics
The spectropolarimetry enabled searches for conventional and generalised Faraday rotation, and spectral depolarisation. The searches showed that most of the detected FRBs had low rotation measures (RMs) $\ll 10^{2}$ rad m-2 Mannings et al. (2023), in contrast to the first repeating FRB which has a high RM Michilli et al. (2018). Similarly, constraints on spectral depolarisation suggest that the scattering media foreground to the ASKAP FRB sample are less magnetoionically active compared to that of repeating FRB sources Uttarkar et al. (2024); Ryder et al. (2023).
@@ 7.5 Intergalactic and circumgalactic media
Our discoveries have also enabled us to uniquely probe circumgalactic media, the similarly diffuse media in galaxy haloes that is difficult to study, but which plays an important role in galaxy formation and evolution. The second FRB we localised (FRB 20181112A) was observed to pass through the halo of an intervening galaxy (Prochaska et al., 2019). Using the resultant limits on pulse broadening and Faraday rotation, we constrained the turbulence and magnetisation in this galaxy halo. Simha et al. (2020) investigated in detail the foreground galaxies towards FRB 20190608B, and Simha et al. (2021) undertook a similar analysis towards FRB 20180924B. We have also been working with the FLIMFLAM collaboration (Lee et al., 2022) to detect galaxy halos along the sightlines to FRBs, with first results on ASKAP FRBs presented in Khrykin et al. (2024).
@@ 8 FUTURE ASKAP FRB SURVEYS
The incoherent searches were a computationally cheap extension to existing ASKAP data-recording systems intended for radio-interferometric imaging on $>10$ s time scales. The searches could be conducted on a single GPU, but at the cost of reduced sensitivity relative to array coherent searches by factor $\approx\sqrt{N_{\rm ant}}$. This reduction in sensitivity provides the opportunity to undertake searches with higher yields with ASKAP by commissioning a new FRB detector. Assuming FRBs are a non-evolving population in a Euclidean Universe, we would expect the detection rate to increase by a factor of $N_{\rm ant}^{3/4}\approx 10$. Coherent surveys are also likely to be able to access a population of FRBs at higher redshift. A coherent FRB search system (the CRAFT Coherent backend, CRACO) is in the final stages of being scientifically commissioned. The system has started detecting FRBs in offline and real-time searches. Additionally, coherent image plane searches are potentially more sensitive to long duration transients, which decorrelate interference and antenna-based gain variations. New classes of long-duration transients of both Galactic Hurley-Walker et al. (2022); Dobie et al. (2024) and extraGalactic origin The CHIME/FRB Collaboration et al. (2022) have been identified. However, the coherent search system will have a smaller field of view so will not be sensitive to FRBs in the outskirts of the primary beam or the side lobes. This will affect FRB detection rates in the outer beams. Even when it is operational we expect to continue with the incoherent sum searches to cross-validate both detectors. The CRACO detector will share the current voltage download system with the ICS detector, delivering high precision astrometry and high time resolution spectro-polarimetry.
@@ Appendix A FRB properties
Table 1 lists the key properties of the FRBs discovered in the ICS survey. Table 2 lists FRB astrometry, derived interferometrically where possible and using the multi-beam localisation method Bannister et al. (2017). For most of the interferometrically localised FRBs we have reported updated positions derived using the CELEBI pipeline Scott et al. (2023). For the the earliest-discovered FRBs (2018094B, 20181112A, 20190102C, 20190608B, 20190611B, and 20190714A), we use positions reported in Day et al. (2021). These FRBs had dedicated astrometric campaigns with the Australian Telescope Compact Array.
@@ Appendix B FRB dynamic spectra
Figures 19 to 23 show the dedispersed dynamic spectra of the FRBs discovered in the ICS survey. The bursts have been arranged in order of increasing dispersion measure. The FRBs are labelled by their internal names.
@@ Appendix C Host galaxy associations and optical photometry
Tables 3 to 6 present the PATH Aggarwal et al. (2021) probabilities of host-galaxy associations for ASKAP-localised FRBs.
@@
Table 3: FRB PATH Associations. For each FRB we list the positions of nearby galaxies, their offset $\theta$ from the burst, angular size $\phi$, and magnitude. Using this information we have calculated PATH probabilities $P(O)$and $P(O|x)$. We note that the automated photometry in the PATH analysis differs from final host galaxy photometry presented in Table 7.
@@
229.7064 $12.3766$ 0.51 0.69 20.68 0.9639 1.0000
229.7087 $12.3777$ 8.83 0.33 24.43 0.0361 0.0000
Table 4: FRB PATH Associations (continued)
RAcand Deccand $\theta$ $\phi$ mag $P(O)$ $P(O|x)$
(deg) (deg) (′′) (′′)
@@
204.5625 $-31.3816$ 4.82 0.64 24.46 0.0205 0.0000
204.5611 $-31.3785$ 7.75 0.58 24.47 0.0204 0.0000
204.5652 $-31.3791$ 9.39 0.66 22.52 0.1014 0.0000
204.5628 $-31.3822$ 7.00 0.25 25.16 0.0121 0.0000
204.5648 $-31.3805$ 7.13 0.27 26.11 0.0062 0.0000
204.5634 $-31.3823$ 7.74 0.28 26.72 0.0041 0.0000
204.5602 $-31.3783$ 9.96 0.28 25.30 0.0110 0.0000
@@
208.8022 $22.4677$ 6.69 0.35 26.09 0.0143 0.0000
208.8048 $22.4666$ 5.02 0.28 26.18 0.0135 0.0000
208.8049 $22.4673$ 6.29 0.37 25.58 0.0204 0.0000
208.8056 $22.4652$ 8.17 0.25 26.95 0.0081 0.0000
208.8057 $22.4672$ 8.69 0.65 22.54 0.2281 0.0000
208.8017 $22.4646$ 7.99 0.47 24.94 0.0324 0.0000
208.8057 $22.4661$ 7.94 0.32 25.36 0.0239 0.0000
208.8013 $22.4649$ 8.54 0.56 24.21 0.0567 0.0000
208.8010 $22.4653$ 8.36 0.38 24.91 0.0332 0.0000
Table 5: FRB PATH Associations (continued)
RAcand Deccand $\theta$ $\phi$ mag $P(O)$ $P(O|x)$
(deg) (deg) (′′) (′′)
@@
303.1165 $-55.3567$ 2.59 0.26 26.70 0.0043 0.0000
303.1177 $-55.3559$ 4.62 0.41 23.74 0.0374 0.0000
303.1111 $-55.3562$ 9.11 0.96 20.35 0.8053 0.0000
303.1185 $-55.3585$ 9.97 0.56 23.18 0.0594 0.0000
303.1174 $-55.3584$ 8.49 0.27 26.13 0.0063 0.0000
Table 6: FRB PATH Associations (continued)
RAcand Deccand $\theta$ $\phi$ mag $P(O)$ $P(O|x)$
(deg) (deg) (′′) (′′)
FRB20230731A: P(U)=0.0, P(U—x)=0.0000
174.6026 $-56.7989$ 2.30 0.28 20.62 0.0833 1.0000
174.5999 $-56.7980$ 5.01 0.29 20.57 0.0878 0.0000
174.5994 $-56.8002$ 5.85 0.39 19.88 0.1751 0.0000
174.6019 $-56.8006$ 5.44 0.29 20.67 0.0793 0.0000
174.5974 $-56.7989$ 7.98 0.68 20.15 0.1331 0.0000
174.5974 $-56.7984$ 8.38 0.64 20.08 0.1427 0.0000
174.6059 $-56.7992$ 8.75 0.38 19.84 0.1820 0.0000
174.6025 $-56.8014$ 8.83 0.29 20.65 0.0806 0.0000
174.6059 $-56.7987$ 8.93 0.22 21.49 0.0361 0.0000

### v2

@@ 7.1 FRB host galaxies
Over the course of the survey, we undertook a series of studies investigating the properties of FRB hosts. Bhandari et al. (2020b) and Heintz et al. (2020) conducted the first demographic studies of FRB host galaxies, finding them to be largely late-type star forming galaxies. Bhandari et al. (2022) extended the analysis to include a larger sample of FRBs and compared the properties of the hosts of apparently one-off FRBs with those of repeating sources. The sample size was insufficient to distinguish any difference between the repeating and non-repeating sources. Based on the Kolmogorov Smirnov tests reported in Bhandari et al. (2022) (which for many had $P_{\rm KS}\approx 0.25$) , which included a sample of $10$ apparently one-off FRBs and $6$ repeating FRBs, the expected sample size needed to be approximately a factor of approximately two larger for each sub-population to reach $95\%$ confidence and a factor of approximately four larger in each to reach $99.9\%$ confidence assuming the distributions had identical distributions.
Gordon et al. (2023) undertook detailed modelling of host galaxy spectral energy distributions to assess the star formation histories of nearly all our FRB hosts identified prior to the beginning of 2022. From these studies it is now clear that the host galaxies of one-off bursts are markedly different from that of the first repeating FRB 20121102A which was found to originate in a metal-poor dwarf galaxy Tendulkar et al. (2017).
@@ 7.2 Burst-emission physics
The spectropolarimetry enabled searches for conventional and generalised Faraday rotation Bera et al. (2024), and spectral depolarisation. The searches showed that most of the detected FRBs had low rotation measures (RMs) $\ll 10^{2}$ rad m-2 Mannings et al. (2023), in contrast to the first repeating FRB which has a high RM Michilli et al. (2018). Similarly, constraints on spectral depolarisation suggest that the scattering media foreground to the ASKAP FRB sample are less magnetoionically active compared to that of repeating FRB sources Uttarkar et al. (2024); Ryder et al. (2023).
@@ 7.5 Intergalactic and circumgalactic media
Our discoveries have also enabled us to uniquely probe circumgalactic media, the similarly diffuse media in galaxy haloes that is difficult to study, but which plays an important role in galaxy formation and evolution. The second FRB we localised (FRB 20181112A) was observed to pass through the halo of an intervening galaxy (Prochaska et al., 2019). Using the resultant limits on pulse broadening and Faraday rotation, we constrained the turbulence and magnetisation in this galaxy halo. Simha et al. (2020) investigated in detail the foreground galaxies towards FRB 20190608B, and Simha et al. (2021) undertook a similar analysis towards FRB 20180924B. We have also been working with the FLIMFLAM collaboration (Lee et al., 2022) to detect galaxy halos along the sightlines to FRBs, with first results on ASKAP FRBs presented in Khrykin et al. (2024) and Huang et al. (2024).
@@ 8 FUTURE ASKAP FRB SURVEYS
The incoherent searches were a computationally cheap extension to existing ASKAP data-recording systems intended for radio-interferometric imaging on $>10$ s time scales. The searches could be conducted on a single GPU, but at the cost of reduced sensitivity relative to array coherent searches by factor $\approx\sqrt{N_{\rm ant}}$. This reduction in sensitivity provides the opportunity to undertake searches with higher yields with ASKAP by commissioning a new FRB detector. Assuming FRBs are a non-evolving population in a Euclidean Universe, we would expect the detection rate to increase by a factor of $N_{\rm ant}^{3/4}\approx 10$. Coherent surveys are also likely to be able to access a population of FRBs at higher redshift. A coherent FRB search system (the CRAFT Coherent backend, CRACO Wang et al., 2024) is in the final stages of being scientifically commissioned. The system has started detecting FRBs in offline and real-time searches. Additionally, coherent image plane searches are potentially more sensitive to long duration transients, which decorrelate interference and antenna-based gain variations. New classes of long-duration transients of both Galactic Hurley-Walker et al. (2022); Dobie et al. (2024) and extra-Galactic origin The CHIME/FRB Collaboration et al. (2022) have been identified. However, the coherent search system will have a smaller field of view so will not be sensitive to FRBs in the outskirts of the primary beam or the side lobes. This will affect FRB detection rates in the outer beams. Even when it is operational we expect to continue with the incoherent sum searches to cross-validate both detectors. The CRACO detector will share the current voltage download system with the ICS detector, delivering high precision astrometry and high time resolution spectro-polarimetry.
@@ Appendix A FRB properties
Table 1 lists the key properties of the FRBs discovered in the ASKAP-CRAFT incoherent sum survey. Table 2 presents FRB astrometry, derived interferometrically where possible, and using the multi-beam localisation method Bannister et al. (2017). For most of the interferometrically localised FRBs we have reported updated positions derived using the CELEBI pipeline Scott et al. (2023). For the the earliest-discovered FRBs (2018094B, 20181112A, 20190102C, 20190608B, 20190611B, and 20190714A), we use positions reported in Day et al. (2021). These FRBs had dedicated astrometric campaigns with the Australian Telescope Compact Array.
@@ Appendix B FRB dynamic spectra
Figures 20 to 24 show the integrated pulse profiles (panel A) and dedispersed dynamic spectra (panel B) of the FRBs discovered in the ICS survey. The bursts have been arranged in order of increasing dispersion measure. FRB designations are listed in the uppper right corner of each plot.
@@ Appendix C Host galaxy associations and optical photometry
Tables 3 to LABEL:tab:path4 present the PATH Aggarwal et al. (2021) probabilities of host-galaxy associations for ASKAP-localised FRBs.
@@
Table 3: FRB PATH Associations. For each FRB we list the positions of nearby galaxies, their offset $\theta$ from the burst, angular size $\phi$, and magnitude. Using this information we have calculated PATH probabilities $P(O)$and $P(O|x)$. We note that the automated photometry in the PATH analysis differs from final host galaxy photometry presented in Table 5. We only report galaxies with $P(O|x)$$>10^{-4}$.
@@
229.7064 $12.3766$ 0.37 0.72 21.18 0.9389 1.0000
@@
Table 4: FRB PATH Associations (continued)
RAcand Deccand $\theta$ $\phi$ mag $P(O)$ $P(O|x)$
(deg) (deg) (′′) (′′)